OpenStudy (anonymous):

It takes Jessie's dog about 13 days to finish a 25-pound bag of dog food. Which is the best estimate of the number of pounds of dog food Jessie's dog eats in one year

Directrix (directrix):

13 days is to 25 pounds of food as a year (365 days) is to x where x is the number of pounds the dog eats in one year. This proportion assumes that the dog's appetite remains constant for the entire year. 13/25 = 365/x 13x = 9125 x = ? @kiarawesson